# Learning Koopman Dynamics for Safe Legged Locomotion with Reinforcement Learning-based Controller

### v1

@@ Abstract
Learning-based algorithms have demonstrated impressive performance in agile locomotion of legged robots. However, learned policies are often complex and opaque due to the black-box nature of learning algorithms, which hinders predictability and precludes guarantees on performance or safety. In this work, we develop a novel safe navigation framework that combines Koopman operators and model-predictive control (MPC) frameworks. Our method adopts Koopman operator theory to learn the linear evolution of dynamics of the underlying locomotion policy, which can be effectively learned with Dynamic Mode Decomposition (DMD). Given that our learned model is linear, we can readily leverage the standard MPC algorithm. Our framework is easy to implement with less prior knowledge because it does not require access to the underlying dynamical systems or control-theoretic techniques. We demonstrate that the learned linear dynamics can better predict the trajectories of legged robots than baselines. In addition, we showcase that the proposed navigation framework can achieve better safety with less collisions in challenging and dense environments with narrow passages.
@@ I INTRODUCTION
Recent advances in reinforcement learning have led to significant improvements in robust and agile quadrupedal locomotion [1, 2, 3, 4, 5, 6]. Extensive experiments have demonstrated robust performance when traversing complex terrains [2, 3, 5], with the potential for efficient training fueled by large computational resources [1] or motion priors [4, 5, 6]. However, learned locomotion policies are often limited to tracking the given velocity command, leaving navigation decisions to a high-level policy.
Some learning-based frameworks also tackle navigation by either training a dedicated high-level policy [7, 8, 9] or developing an end-to-end policy using diverse navigation maps and additional navigation rewards [10, 11, 12]. While these approaches are effective, they demand significant computational resources and extensive user expertise for trail-and-error-based parameter tuning. Moreover, learning-based approaches often demonstrate degraded performance in unseen environments. As such, these approaches lack strong safety guarantees despite appearing to be safe under specific experimental conditions.
Unlike learning-based approaches, conventional model-based approaches tend to integrate obstacle avoidance constraints with model predictive control (MPC) for safe robot navigation. However, these approaches assume explicit knowledge of the robot’s underlying dynamical model. This model tends to be highly complex due to the hybrid dynamics of frequent making and breaking of contacts [13], and highly nonlinear whole-body dynamics [14]. To circumvent such complexity and facilitate constrained optimization, existing methods resort to model simplifications [13, 15, 16], but sacrifice accuracy in the process.
We contribute an easy-to-use framework (Fig. 1 for an overview) for safe navigation that can be used in conjunction with any locomotion controller, learned or otherwise. Our key idea is to approximately encode the closed-loop dynamics of legged motion using a unified linear model that captures the co-evolution of state variables (e.g., pose and velocity) over time. In particular, we use Dynamic Mode Decomposition (DMD) [17] to efficiently learn unified linear models from data. We also propose the use of approximated Koopman operator [18] in conjunction with DMD to encode underlying nonlinear dynamics using a linear system defined in a higher-dimensional space.
Exploiting the fact our learned dynamics are linear, we employ a conventional model-predictive controller (MPC) to generate the desired velocity commands for collision-free navigation in dense and narrow environments. Because the dynamic constraints are linear, our framework is simple and can be used with modest computational resources. We do not assume any knowledge of the underlying dynamics or its structure, and our framework can learn from offline locomotion data (e.g., rollouts of pre-trained controllers with random control commands).
Figure 1: Our approach learns to approximately represent the high-level closed-loop dynamics of a given learning-based locomotion controller using a unified linear model, which we then utilize to ensure model-based safety.
@@
We conducted extensive experiments to evaluate the effectiveness of our framework. First, we compared the long-term prediction accuracy of our unified linear and koopman-based models against that of a component-based linear model [19], and a simple linear integrator model. Our results indicate that both our models consistently outperform the baselines in terms of long-term prediction accuracy on the withheld dataset, with our Koopman-based model performing better than our unified linear model.
Second, we compared each model’s ability to enable safe navigation when used in conjunction with MPC (See Fig. 2 for examples). Specifically, we conducted our evaluation on three different maps of varying difficulty. Our results suggest that our models can enable safe navigation across all the maps, outperforming the component-wise linear and the simple integrator models.
The paper is organized as follows: In Sec. II, we discuss related works and contextualize our contributions. In Sec. III, we introduce DMD and Koopman Operator theory. In Sec. IV, we detail our framework, including offline data collection, model learning, and model-predictive safety. In Sec. V, we discuss experimental results. In Sec. VI, we summarize our contributions and suggest future directions.
@@ II-A Safe Legged Locomotion
It is important to ensure safety during legged robots’ dynamic motions with underactuated bases. Previously, safety has been widely discussed within the context of model-based control [20, 14]. Recent advances in learning-based locomotion [21, 1] frameworks have demonstrated impressive robustness, but there are only limited works that discuss the interpretability [22] and guarantee of safety [23]. A common approach in this context is to decouple a low-level locomotion controller and a high-level navigation policy, where a high-level policy adopts conventional methods such as model predictive control and control barrier functions [24, 25, 26]. Li et al. [19] propose a method that leverages low-dimensional linear models identified from bipedal locomotion policies trained via reinforcement learning. Liao et al. [27] models robots and obstacles as polytopes [28] to enable narrow space navigation. However, they require access to a low-level whole-body controller, which is not available in many real-world scenarios. Instead of using analytic formulations, there exist works to model low-level behaviors using neural networks [29, 30]. However, the learned dynamics are often highly nonlinear, which makes it infeasible to adopt conventional optimization methods and requires additional techniques, such as an informed trajectory sampler [30].
@@ II-B Koopman Operator for Robotics
The Koopman operator theory [31], which offers a linear perspective on nonlinear dynamical systems, has gained growing interest in robotics for its potential to simplify complex dynamical systems [32] and high sample efficiency [33]. They have been successfully utilized in various domains such as spherical robots [32], drones [34, 35, 36], and dexterous manipulation [33, 37, 38]. Leveraging their linearity, Koopman operator methods have widely been combined with model predictive controls to obtain desired command tracking [32, 39], enforce better stability, or obtain safety. The Koopman operator has also proven useful in learning the reference dynamics from trajectory data, which demonstrates superior sample efficiency compared to deep neural networks in dexterous manipulation [33]. In this work, we learn the high-level dynamics of the velocity-tracking policy and leverage a model predictive control framework to obtain safe commands to navigate within environments with obstacles.
III Preliminaries: DMD and Koopman Operators
We begin by providing a brief introduction to DMD and Koopman operator theory [18].
@@ III Preliminaries: DMD and Koopman Operators
Koopman Representation: Consider a discrete-time autonomous nonlinear dynamical system $\mathrm{x}_{t+1}=F(\mathrm{x}_{t}),$ where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state at time $t$, and $F(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a nonlinear transition function. To precisely represent this nonlinear dynamical system as a linear system, we introduce a set of observables using the so-called lifting function $\psi:\mathcal{X}\rightarrow\mathcal{O}$, where $\mathcal{O}$ is the space of observables. We can now define the Koopman Operator $\mathcal{K}$, an infinite-dimensional operator on the lifting function $\psi(\cdot)$ for the discrete-time nonlinear system as follows
$[\mathcal{K}\psi]=\psi(F(\mathrm{x}_{t}))=\psi(\mathrm{x}_{t+1})$ (1)
@@
If the observables belong to a vector space, $\mathcal{K}$ can be seen as an infinite-dimensional linear map that describes the evolution of the observables: $\psi(\mathrm{x}_{t+1})=\mathcal{K}\psi(\mathrm{x}_{t})$.
@@
In practice, we do not benefit from this representation since it is infinite-dimensional. However, we can approximate $\mathcal{K}$ using a matrix $K\in\mathbb{R}^{p\times p}$ and define a finite set of observables $\psi(t)\in\mathbb{R}^{p}$. Thus, we can rewrite the relationship as $\psi(\mathrm{x}_{t+1})=K\psi(\mathrm{x}_{t})+r(\mathrm{x}_{t}),$ where $r(\mathrm{x}_{t})\in\mathbb{R}^{p}$ is the residual error caused by the finite dimensional approximation, which can be arbitrarily reduced based on the choice of the lifting function, $p$ and $\psi(\cdot)$.
Further, let’s consider a controlled dynamical system $\mathrm{x}_{t+1}=F(\mathrm{x}_{t},\mathrm{u}_{t}),$ where $\mathrm{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$. Now we aim to satisfy $\psi(\mathrm{x}_{t+1},\mathrm{u}_{t+1})=\mathcal{K}\psi(\mathrm{x}_{t},\mathrm{u}_{t})$. Here, a common strategy [32] is to make $\psi(\cdot)$ concatenate the lifted states with the control commands: $\psi(\mathrm{x}_{t},\mathrm{u}_{t})=[\phi(\mathrm{x}_{t});\mathrm{u}_{t}],$ where $\phi(\cdot)$ is the state-dependent lifting function. Finally, we can obtain the new finite approximation as $[\phi(\mathrm{x}_{t+1});\mathrm{u}_{t+1}]=K[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]+r(\mathrm{x}_{t},\mathrm{u}_{t}),$ where $K=\begin{bmatrix}A\in\mathbb{R}^{p\times p}&B\in\mathbb{R}^{p\times m}\\ C\in\mathbb{R}^{m\times p}&D\in\mathbb{R}^{m\times m}\end{bmatrix}$.
@@
Learning Koopman Operator from Data via DMD: The matrix operators $K$ can be inferred from a dataset $\mathcal{D}=[(\mathrm{x}_{1},\mathrm{u}_{1}),\cdots,(\mathrm{x}_{T},\mathrm{u}_{T})]$, which contains the system evolutions and control inputs. Given the choice of observables $\phi(\cdot)$, $K$ is computed by minimizing the residual errors. Specifically, we can obtain $K$ from $\mathcal{D}$ by minimizing the cost function $J(K)$ given below:
$\footnotesize J(K)=\frac{1}{2}\sum_{t=1}^{t=T-1}\|r(\mathrm{x}_{t},\mathrm{u}_{t})\|^{2}=\frac{1}{2}\sum_{t=1}^{t=T-1}\|[\phi(\mathrm{x}_{t+1};\mathrm{u}_{t+1})]-K[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]\|^{2}.$ (2)
Note minimizing $J(K)$ amounts to solving a least-square problem, whose analytical solution can be obtained using the DMD technique [17, 40]: $K=XY^{\dagger}$, where $X=\frac{1}{T-1}\sum_{t=1}^{t=T-1}[\phi(\mathrm{x}_{t+1});\mathrm{u}_{t+1}]\otimes[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]$, and $Y=\frac{1}{T-1}\sum_{t=1}^{t=T-1}[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]\otimes[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]$. Here, $Y^{\dagger}$ denotes the Moore–Penrose inverse of $Y$, and $\otimes$ denotes the outer product. In the standard application of state prediction without predicting the flow of $\mathrm{u}_{t}$, after obtaining the matrix $K$, we can use $\phi(\mathrm{x}_{t+1})=A\phi(\mathrm{x}_{t})+B\mathrm{u}_{t}$ to model the system evolution under control inputs, which facilitates the linear control synthesis [41, 42, 35, 43, 44].
@@ IV Safe Navigation with Learned Koopman Forward Dynamics
This section presents a novel safe navigation framework that leverages the learned Koopman dynamics. A typical model-predictive control (MPC) has been widely adopted as a theoretical backbone for safe navigation. However, the high-level forward dynamics of a legged robot are hard to analyze because it is a closed-loop system with a nonlinear control policy and hybrid contact dynamic, which prevents the existing safety-aware control algorithms from capturing the legged locomotion behaviors accurately.
Instead, we capture a legged robot’s high-level forward dynamics as a linear evolution by adopting the Koopman operator theory, which projects its state space into a high-dimensional feature space. Due to its linearity, we can easily combine the learned dynamics with the existing MPC-based safe navigation frameworks.
@@ IV-A Low-level locomotion controller
Our framework assumes a velocity-tracking controller that takes as input the target velocity and outputs motor commands. Our framework is agnostic to the implementation of a low-level controller. We can use either model-based control theory [14] or a learning-based algorithm, which works best for the given scenario. In our implementation, we train a controller using Proximal Policy Optimization (PPO) inspired by the work of Rudin et al. [1]. We train Unitree’s Aliengo robot using an IsaacGym simulator with 4096 parallel environments for 30 minutes in NVIDIA RTX 4090 GPU until the training converges.

### v2

# Safe Navigation of Bipedal Robots via Koopman Operator-Based Model Predictive Control

@@ Abstract
Nonlinearity in dynamics has long been a major challenge in robotics, often causing significant performance degradation in existing control algorithms. For example, the navigation of bipedal robots can exhibit nonlinear behaviors even under simple velocity commands, as their actual dynamics are governed by complex whole-body movements and discrete contacts. In this work, we propose a novel safe navigation framework inspired by Koopman operator theory. We first train a low-level locomotion policy using deep reinforcement learning, and then capture its low-frequency, base-level dynamics by learning linearized dynamics in a high-dimensional lifted space using Dynamic Mode Decomposition. Then, our model-predictive controller (MPC) efficiently optimizes control signals via standard quadratic objective and the linear dynamics constraint in the lifted space. We demonstrate that the Koopman-based model more accurately predicts bipedal robot trajectories than baseline approaches. Furthermore, we show that the proposed navigation framework achieves improved safety with better success rates in dense environments with narrow passages.
@@ I INTRODUCTION
Nonlinear dynamics have long been a fundamental challenge in robotics, which often leads to significant difficulties in modeling, prediction, and control. Legged robot navigation is one of the examples because even simple velocity commands can induce highly nonlinear behaviors due to its complex hybrid dynamics. Researchers have explored both model-based and learning-based approaches. Model-based methods often integrate obstacle avoidance constraints with model predictive control (MPC), leveraging explicit dynamic models for safe navigation [1, 2, 3, 4]. On the other hand, learning-based frameworks have achieved impressive results in quadrupedal locomotion and navigation, either by training high-level policies [5, 6, 7] or developing end-to-end policies [8, 9, 10]. While both lines of research have been successful, model-based approaches are constrained by the complexity of accurate dynamics modeling and the increased computation cost, whereas learning-based methods demand extensive training data and tend to show degraded performance in unseen environments.
Researchers have explored various algorithms that combine the strengths of both model-based and learning-based approaches [11, 12]. A notable example is the work of Li et al. [13], which identifies low-dimensional linear models of RL-based closed-loop dynamics in legged robots via system identification. Their results demonstrate that the learned models can successfully enable safety-critical optimal control, such as MPC with control barrier functions, on physical systems. However, this method learns low-dimensional, component-wise dynamics, which may not sufficiently capture the nonlinearity of high-level behaviors in more challenging scenarios.
Figure 1: Our approach first trains a locomotion controller on a bipedal robot and then learns the linearized high-level dynamics using Koopman operator theory. Once the dynamics are learned, we use it to develop an MPC-style safe navigation controller.
In this work, we leverage Koopman operator theory to represent the nonlinear dynamics of bipedal robots as linear transitions in a high-dimensional lifted space [14, 15]. We first train a low-level locomotion policy via deep reinforcement learning and then capture its low-frequency, base-level dynamics using Dynamic Mode Decomposition (DMD). Unlike existing approaches that rely on simplified analytic models or black-box neural networks, our framework provides a simple yet accurate representation of dynamics. Building on this model, we formulate an MPC framework using a standard quadratic objective and a linear dynamics constraint, enabling both computational efficiency and reliable performance for safety-critical navigation tasks.
@@
We conducted experiments to evaluate the effectiveness of our framework. In terms of long-term prediction accuracy, both Koopman-based models consistently outperformed a component-wise linear model [13], a simple linear integrator, and a unified linear model, with the Koopman-based model achieving the highest accuracy. We further assessed each model’s ability to support safe navigation of a G1 humanoid when combined with MPC across four maps of varying difficulty (Fig. 2). Our results demonstrate that the Koopman model enabled reliable safe navigation by achieving higher success rates compared to the baselines. The main contributions of this work are:
  • We propose a novel Koopman-based safe navigation control framework for bipedal robots, which leverages learned linear Koopman dynamics and model-predictive control.
  • We provide a comprehensive evaluation of various forward dynamics models on bipedal robots, including an integrator, component-wise linear model, a unified linear model, and Koopman dynamics models with different lifting functions.
  • We demonstrate improved safety of the proposed framework compared to the baselines.
@@ II-A Safe Legged Locomotion
It is important to ensure safety during legged robots’ dynamic motions with underactuated bases. Previously, safety has been widely discussed within the context of model-based control [16, 2]. Recent advances in learning-based locomotion frameworks [17, 18] have demonstrated impressive robustness, but there are only limited works that discuss the interpretability [19] and guarantee of safety [20]. A common approach in this context is to decouple a low-level locomotion controller and a high-level navigation policy, where a high-level policy adopts conventional methods such as model predictive control and control barrier functions [21, 22, 23]. Li et al. [13] propose a method that leverages low-dimensional linear models identified from bipedal locomotion policies trained via reinforcement learning. Liao et al. [24] models robots and obstacles as polytopes [25] to enable narrow space navigation. However, they require access to a low-level whole-body controller, which is not available in many real-world scenarios. Instead of using analytic formulations, there exist works to model low-level behaviors using neural networks [26, 27]. However, the learned dynamics are often highly nonlinear, which makes it infeasible to adopt conventional optimization methods and requires additional techniques, such as an informed trajectory sampler [27].
@@ II-B Koopman Operator for Robotics
The Koopman operator theory [28], which offers a linear perspective on nonlinear dynamical systems, has gained growing interest in robotics for its potential to simplify complex dynamical systems [29] and high sample efficiency [30]. They have been successfully utilized in various domains such as spherical robots [29], drones [31, 32, 33], and dexterous manipulation [30, 34, 35]. Leveraging their linearity, Koopman operator methods have widely been combined with model predictive control to obtain desired command tracking [29, 36], enforce better stability, or improve safety. The Koopman operator has also proven useful in learning the reference dynamics from trajectory data, which demonstrates superior sample efficiency compared to deep neural networks in dexterous manipulation [30]. In this work, we learn the high-level dynamics of the velocity-tracking policy and leverage a model predictive control framework to obtain safe commands to navigate within environments with obstacles.
III Preliminaries: Learning Koopman Dynamics
This section introduces Koopman operator theory, a framework that enables representing nonlinear dynamics through linear evolution in a lifted space [15].
III-1 Koopman Representation
@@
Consider a nonlinear system $\mathrm{x}_{t+1}=F(\mathrm{x}_{t}),$ where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ denotes the state at time $t$, and $F(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a nonlinear transition function. We introduce a set of observables using the so-called lifting function $\psi:\mathcal{X}\rightarrow\mathcal{O}$, where $\mathcal{O}$ denotes the space of observables, to capture nonlinear dynamics as linear relationships between observables. We define the Koopman operator $\mathcal{K}$ acting on an observable function $\psi$ as:
$(\mathcal{K}\psi)(x_{t})=\psi(F(x_{t}))=\psi(x_{t+1}).$
@@
If the observables form a vector space, $\mathcal{K}$ can be regarded as an infinite-dimensional linear operator that governs their evolution: $\psi(\mathrm{x}_{t+1})=\mathcal{K}\psi(\mathrm{x}_{t})$.
@@
In practice, this representation is not directly useful because it is infinite-dimensional, which makes computation infeasible. To address this limitation, we approximate $\mathcal{K}$ with a finite-dimensional matrix $K\in\mathbb{R}^{p\times p}$ and define a corresponding finite set of observables $\psi(t)\in\mathbb{R}^{p}$. The system can then be represented as $\psi(\mathrm{x}_{t+1})=K\psi(\mathrm{x}_{t})+r(\mathrm{x}_{t}),$ where $r(\mathrm{x}_{t})\in\mathbb{R}^{p}$ is the residual error caused by the finite-dimensional approximation. This error can be reduced by appropriately selecting the lifting function, the dimension $p$, and the mapping $\psi(\cdot)$.
In addition, we can illustrate a controlled dynamical system $\mathrm{x}_{t+1}=F(\mathrm{x}_{t},\mathrm{u}_{t}),$ with a control $\mathrm{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{m}$, which yields $\psi(\mathrm{x}_{t+1},\mathrm{u}_{t+1})=\mathcal{K}\psi(\mathrm{x}_{t},\mathrm{u}_{t})$. A common strategy [29] is to design $\psi(\cdot)$ by lifting states with a state-specific function $\phi(\cdot)$ and concatenating them with the control signal $\psi(\mathrm{x}_{t},\mathrm{u}_{t})=[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]$, which yields:
$\phi(\mathrm{x}_{t+1});\mathrm{u}_{t+1}]=K[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]+r(\mathrm{x}_{t},\mathrm{u}_{t}).$
If we skip control prediction, we obtain:
$\phi(\mathrm{x}_{t+1})=A\phi(\mathrm{x}_{t})+B\mathrm{u}_{t}+r(\mathrm{x}_{t},\mathrm{u}_{t}).$
III-2 Learning Koopman Dynamics
@@
The matrix operators $K$ can be learned from a dataset $\mathcal{D}=[(\mathrm{x}_{1},\mathrm{u}_{1}),\cdots,(\mathrm{x}_{T},\mathrm{u}_{T})]$, which contains the state trajectories and control inputs. Given the choice of state observables $\phi(\cdot)$, $K$ can be computed by minimizing the cost function $J$:
$\footnotesize J(K)=\sum_{t=1}^{t=T-1}\|r(\mathrm{x}_{t},\mathrm{u}_{t})\|^{2}=\sum_{t=1}^{t=T-1}\|[\phi(\mathrm{x}_{t+1});\mathrm{u}_{t+1}]-K[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]\|^{2}.$ (1)
Because this is a least-square problem, we can simply obtain an analytical solution. We can define $X$ and $Y$ as:
$X$ $=\frac{1}{T-1}\sum_{t=1}^{t=T-1}[\phi(\mathrm{x}_{t+1});\mathrm{u}_{t+1}]\otimes[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]$
$Y$ $=\frac{1}{T-1}\sum_{t=1}^{t=T-1}[\phi(\mathrm{x}_{t});\mathrm{u}_{t}]\otimes[\phi(\mathrm{x}_{t});\mathrm{u}_{t}],$
where $\otimes$ denotes the outer product. Then, we can take the Moore-Penrose inverse $K=XY^{\dagger}$ as suggested in [14, 37].
@@ IV Safe Navigation with Learned Koopman Forward Dynamics
This section presents a novel safe navigation framework that leverages the learned Koopman dynamics. Model-predictive control (MPC) has been widely adopted as a theoretical backbone for safe navigation. However, the high-level forward dynamics of a bipedal robot are difficult to analyze because it is a closed-loop system with a nonlinear control policy and hybrid contact dynamics, which can degrade the performance of the existing safety-aware control algorithms.
Instead, we capture a robot’s high-level forward dynamics as a linear evolution of high-dimensional features by adopting the Koopman operator theory. Due to its linearity, this enables effective integration with the existing MPC-based safe navigation algorithms.
@@ IV-A Low-level Locomotion Controller
Our framework assumes a velocity-tracking controller that takes as input the target velocity and outputs motor commands. While high-level dynamics learning can be agnostic to the implementation of a low-level controller, we opt for deep RL that can learn a robust control policy from a relatively simple reward function.
Following the reward structure from [38], we train a locomotion controller for the Unitree G1 robot [39] using Proximal Policy Optimization (PPO) [40]. The low-level control policy $\pi_{\theta}(a_{t}\mid o_{t})$ maps proprioceptive observations $o_{t}$ to actions $a_{t}\in\mathbb{R}^{12}$, which correspond to target joint angles for the 12 lower body joint actuators. A PD controller then converts these targets into torques using joint-specific stiffness and damping gains. The policy observes $o_{t}\in\mathbb{R}^{47}$:
$o_{t}=[{\omega}_{t},g_{t},\hat{\mathbf{v}},q,\dot{q}_{t},a_{t-1},\sin(2\pi c_{t}),\cos(2\pi c_{t})]$
The observation vector consists of the base angular velocity $\omega_{t}\in\mathbb{R}^{3}$, the projected gravity vector $g_{t}\in\mathbb{R}^{3}$, and the command linear and angular velocities $\hat{\mathbf{v}}=[\hat{v_{x}},\hat{v_{y}},\hat{\omega}]\in\mathbb{R}^{3}$. It further includes the joint positions and velocities $q,\dot{q}\in\mathbb{R}^{12}$, together with the previous action $a_{t-1}\in\mathbb{R}^{12}$. To aid in learning rhythmic gaits, a 2D gait phase clock $\sin(2\pi c_{t}),\cos(2\pi c_{t})]$ is included, where the phase clock $c_{t}$ cycles with a period of 0.8 seconds. The policy is modeled using a recurrent neural network(LSTM).
Training is performed in the IsaacGym simulator [41] with 4,096 parallel environments for 7,000 steps, requiring approximately two hours on an NVIDIA RTX 4090 GPU.